# On Your Own: Pro-level Autonomous Drone Racing in Uninstrumented Arenas

### v1

@@ Abstract
Drone technology is proliferating in many industries, including agriculture, logistics, defense, infrastructure, and environmental monitoring. Vision-based autonomy is one of its key enablers, particularly for real-world applications. This is essential for operating in novel, unstructured environments where traditional navigation methods may be unavailable. Autonomous drone racing has become the de facto benchmark for such systems. State-of-the-art research has shown that autonomous systems can surpass human-level performance in racing arenas. However, direct applicability to commercial and field operations is still limited as current systems are often trained and evaluated in highly controlled environments. In our contribution, the system’s capabilities are analyzed within a controlled environment—where external tracking is available for ground-truth comparison—but also demonstrated in a challenging, uninstrumented environment—where ground-truth measurements were never available. We show that our approach can match the performance of professional human pilots in both scenarios. We also publicly release the data from the flights carried out by our approach and a world-class human pilot: github.com/tii-racing/drone-racing-dataset
@@ I Introduction
The last two decades of rapid progress in AI have seen machine learning systems challenge, and even defeat, humans in several tasks (from image recognition, to the game of Go, to self-driving). Since the 2019 AlphaPilot AI Drone Innovation Challenge [2, 3], drone racing has emerged as the de facto benchmark for evaluating vision-based aerial autonomy. In 2023, state-of-the-art research has demonstrated that autonomous racing drones can surpass human-level performance in controlled racing arenas [4]. However, the applicability of fast, vision-based autonomous flight to practical commercial operations is still limited, as current systems are typically trained and evaluated in highly controlled (often externally instrumented) environments.
@@
  • Presenting a perception and control stack for autonomous drone racing that does not require to be fine-tuned with ground truth for residual estimation (and has proven to be resilient to multiple lighting conditions).
  • Releasing pro-level piloting data from the instrumented track using the same format as in [5], adding six new flights by a world champion pilot, including 240.77s of total flying time, 2342.98m of traveled distance, at a top speed of 21.29m/s.
@@
This 2025 Quadcopter Research 25.630m/s
5.047s 21.157m/s
5.601s Human
@@ II Related Work
In car racing, since the 2004/2005 DARPA Grand Challenges, various leagues and research challenges have been established, including the Indy Autonomous Challenge (IAC) and the Abu Dhabi Autonomous Racing League (A2RL) [6]—the latter organizing “Man vs Machine” events for both drone and car racing [7]—aimed at accelerating progress in performance, research, and entertainment.
@@ III Drone Platform
The components listed in this section allowed us to create an autonomous drone capable of speeds over 25m/s along an agile indoor trajectory and weighing just 665.5g (without battery).
@@ III-A1 Motors and Propellers
The quadrotor features the T-motor F60 PRO V 2020KV motors, as in [5], paired with HQProp HeadsUp R38 propellers to accommodate for the professional pilot’s unanimous preference for a lower blade pitch. This setup results in a Thrust-to-Weight Ratio (TWR) of $\sim$7 at full battery capacity, measured using a thrust bench.
@@ III-A3 Compute
As companion board, we use the NVIDIA Orin NX module, installed on the A603 carrier board by Seeedstudio. The Orin NX module comes with JetPack 5.1.2 (i.e., Linux Kernel 5.10, Ubuntu 20.04-based root file system, and CUDA 11.4 support). We select the MAXN power mode to maximize core usage and clock frequencies.
@@ III-B1 Autopilot
Our drone requires low-latency, real-time reactive control. This is managed by the FC, running Betaflight (BF) [13] 4.3.2 firmware, and carefully tuned for the platform. The FC is interfaced using the MultiWii Serial Protocol (MSP), enabling efficient bidirectional data exchange between the FC and the onboard computer. This allows for IMU readings at a rate of up to 500Hz over a 1MBaud serial connection (much more than the 10Hz obtainable with the SBUS protocol, as reported in [12]). Additionally, the MSP Override feature provides the onboard computer with complete control over the flight channels. A real-time Linux kernel setup—with priority granted to the MSP process—was crucial to achieve the performance reported in this letter.
@@ III-B2 Vison and Autonomy
Figure 4: Vision stack schematics illustrating the training procedure (top) and the detection pipeline (bottom). The training procedure comprises three steps (designed to minimize human supervision while preserving accuracy): (i) training on the RATM dataset [5]; (ii) auto-labeling, followed by human refinement in CVAT; and (iii) labeling track images with Grounding DINO from human examples. By the final stage, fine-tuning for peak performance in new environments requires as few as 80 human-corrected frames. In the detection pipeline, YOLOv8n [15] detects gate bounding boxes, while MobileNetV3-Small [16] estimates their keypoints. Leveraging GPU acceleration (see Sec. IV-A), the stack achieves per-frame latency under 30 ms on $640\times 640$ images. This yields lower delay and larger input resolution compared to [4], resulting in higher pixel-level precision and potentially improved PnP accuracy.
@@ IV-A Vision Stack
For gate detection, we use You Only Look Once (YOLO), specifically, the YOLOv8n [15] model with $3.2$ million parameters, one class (gates), and input size ($\times640$).
For corner detection, the pixels corresponding to each detected gate region are cropped and forwarded to a custom key-points detection model, MobileNetV3-Small [16] with $1.1$ million parameters. The input size of the model is ($\times256$) and its weights are initially pretrained for a classification task using the ImageNet-1K dataset [17].
@@ IV-B State Estimation Stack
Our state estimation stack leverages VIO from the Intel T265. Although VIO provides highly accurate short-term estimates, its performance degrades over time, leading to significant drift, especially during fast maneuvers that introduce motion blur. To mitigate this drift, we incorporate detections of the only predefined landmarks in drone racing: the gates. By identifying the corners of the gates in the captured images, the drone’s pose relative to the gate is computed by solving the Perspective-n-Point (PnP) problem. The global pose of the drone is then refined using the gate’s known location in the track layout.
Using fixed exposure and carefully tuning the exposure time and gain, we obtain good image brightness with minimal motion blur, ensuring reliable VIO. We enable “mapping”, allowing the device to build and update an internal map for loop closure and small drift correction. In contrast, we disable “relocalization”, which attempts to re-align the estimate after large drifts, and “pose jumping”, which can cause discontinuous estimates. This configuration provided the best balance of VIO accuracy and reliability, while still allowing us to correct larger drifts using gate measurements. The right camera operates at 30Hz and its frames are synchronized with the corresponding VIO estimates.
@@
The filter state and covariance are initialized to zero, while the process noise is set to $\sigma_{a}^{2}=8$. The filter is then updated whenever a new measurement $\mathbf{z}_{k}$ (a pose estimate from a gate detection) is available, using the Kalman filter equations:
@@
in which $K_{k}$ is the Kalman gain, $R$ is the measurement covariance matrix, estimated via Monte-Carlo sampling as described in [4], and $H_{k}$ is the measurement matrix. When several gates are detected in a single camera frame, all relative pose estimates are stacked and processed in the same Kalman filter update step. The resulting drift estimate is constantly used to correct the VIO position estimate. As a final step, the state estimate is refined by fusing it with IMU data from the FC, which operates at 500Hz. This fusion is performed using an Extended Kalman Filter (EKF) with state vector $\mathbf{x}=[\mathbf{q}^{T},\mathbf{p}^{T},\mathbf{v}^{T},\mathbf{b}_{\omega}^{T},\mathbf{b}_{a}^{T}]^{T}\in\mathbb{R}^{16}$ and error state vector $\mathbf{e}=[\mathbf{\psi},\mathbf{\theta},\mathbf{\phi},\mathbf{p}^{T}]^{T}\in\mathbb{R}^{6}$, in which $\mathbf{q}$ represents the drone’s orientation quaternion and $\mathbf{\psi},\mathbf{\theta},\mathbf{\phi}$ its Euler angles representation, $\mathbf{p}$ is its position, $\mathbf{v}$ is its linear velocity, $\mathbf{b}_{\omega}$ is the bias of the gyroscope and $\mathbf{b}_{a}$ is the bias of the accelerometer. The propagation step follows equations:
@@
Where $\mathbf{P}$ is the state covariance matrix, $\mathbf{F}$ is the state transition matrix and $\mathbf{W}$ is the Jacobian of the error-state transition model with respect to process noise. The update step is then defined as in 3, 4, 5, but in this case $\mathbf{R}$ is a static diagonal covariance matrix.
@@ IV-C1 Time Optimal Trajectory Generation
Each gate was assigned two waypoints, positioned at the center of the gate along the y- and z-axes. Along the x-axis, the waypoints were placed at -0.4m and +0.4m (-0.4m and +1.25m for the Split-S). This displacement is defined in gate-frame and it is transformed in world-frame using the known gate yaw angle (in radiant) specified in Fig. 5. The world- and gate-reference frame are shown in Fig. 1.
@@ V Racing Tracks
We used two different tracks, shown in Fig. 5, for our experiments. The first track was set up in a flight arena equipped with a motion capture (MoCap) system, providing ground-truth data to analyze speed, gate-to-gate times, traveled distance, state estimation, and trajectory tracking. The second track was prepared in a larger, uninstrumented hangar. In this case, quantitative evaluation was limited to lap and sector times.
@@ V-A2 Motion Capture
The instrumented arena is equipped with a 32-camera Arqus A12 Qualisys MoCap system [^3] https://www.qualisys.com/cameras/arqus, capable of tracking the 6DoF poses of specified rigid bodies with millimeter-level precision at a frequency of 275Hz. The drone presented in Sec. III-A was outfitted with five 25mm markers, three on 3D-printed arm extensions, and two near the center of the frame.
@@ V-B1 Technical Specifications
In the larger, uninstrumented hangar, we reproduced, as faithfully as possible, the Track Split-S from [4]. We positioned the gates with the support of a total station, resulting in position errors in the order of a few centimeters. The orientations of the gates had errors of at most ten degrees. We should also note that the bottom gate of the Split-S in our track is 20cm lower compared to the original track due to a different gate interlocking system. The gates placement is similar to that of Track RATM. Both tracks feature the Split-S maneuver and similar z-axis variations, although this track has softer curves and lacks the sharp hairpin and spiral segment. The gate poses of the original track are published in the supplementary material of [4].
@@ VI-A Participants
Three professional pilots participated in the experiment, divided into two sessions. The first session included two top-ten-ranked pro-pilots from the Drone Champions League (DCL), namely Thomas Kund (Star23467) and Krutharth M. C. (Ion FPV). The second session featured a world-champion pilot, Minchan Kim (MCK), who had earned multiple titles in recent years—in MultiGP Championship, FAI World Drone Racing Championship (WDRC), and Drone Racing League (DRL)—and carried to victory his team in the Drone Champions League.
@@ VI-B Runs
The pilots were also let free to choose how many batteries or hours of practice to use each day, as well as the number of laps to complete with each battery. The autonomous drone always ran one lap per battery instead.
@@ VI-C Performance
The instrumented evaluation allowed the comprehensive analysis of multiple quantitative metrics including lap times, average speeds, and path lengths. We recorded data for three human pilots and our autonomous system, the latter using both external state estimation (MoCap) as well as fully on-board autonomy (VIO). The results of these flights are presented in Table II and Fig. 6. The autonomous systems (MoCap and VIO) surpassed the humans in average lap time and top speed, although they fell short to one out of the three pilots in terms of best individual results (achieved by MCK). In the instrumented arena, the autonomous system showed a clear reliability advantage, completing all runs without any crash, whereas the human pilots suffered multiple incidents.
@@ VI-D Dataset
Six flights by the fastest pilot, MCK, totaling 240.77s of flying time and 2342.98m of traveled distance, at a top speed of 21.29m/s are available at github.com/tii-racing/drone-racing-dataset, using the same format as in [5].
Figure 7: Comparison of the trajectories followed during the Split-S maneuver by the fastest pilot (MCK) and our autonomous system using only onboard sensing (VIO), sorted by lap time, as recorded by the motion capture system on the instrumented track.
@@
Figure 9: Results from the uninstrumented track showing the lap time distributions and the number of crashes for the fastest pilot (MCK) and our autonomous system.
@@ VII Uninstrumented Results
The scenario was not entirely unknown, as the basic data required for the vision stack were available. To support the vision module, images of the new location were captured and used to fine-tune the corner detection algorithm. This adjustment allowed to accommodate the significant variations in lighting conditions caused by sunlight streaming through the hangar windows (as tests had to be carried out at all hours of the day). The gate map, essential to support the state estimation module, was created using a total station. Although accurate, this method was considerably less precise than the map generated by the MoCap system.
@@ VII-C Performance
MCK achieved both a lower average and best lap time (see Table II, Fig. 9). The autonomous drone, that was 1.27% faster than MCK on average lap time in the instrumented track, gave up a 3.65% margin to the fastest human. Note that the second and third ranked pro-pilots were $\sim$40-60% slower than MCK in the instrumented track. MCK had 2 crashes, whereas the autonomous system crashed 4 times—3 of which were due to collisions with the human pilot, who was able to recover in two of these cases.

### v2

@@ Abstract
Drone technology is proliferating in many industries, including agriculture, logistics, defense, infrastructure, and environmental monitoring. Vision-based autonomy is one of its key enablers, particularly for real-world applications. This is essential for operating in novel, unstructured environments where traditional navigation methods may be unavailable. Autonomous drone racing has become the de facto benchmark for such systems. State-of-the-art research has shown that autonomous systems can surpass human-level performance in racing arenas. However, the direct applicability to commercial and field operations is still limited, as current systems are often trained and evaluated in highly controlled environments. In our contribution, the system’s capabilities are analyzed within a controlled environment—where external tracking is available for ground-truth comparison—but also demonstrated in a challenging, uninstrumented environment—where ground-truth measurements were never available. We show that our approach can match the performance of professional human pilots in both scenarios. We also publicly release the data from flights carried out by a world-class human pilot: github.com/tii-racing/drone-racing-dataset.
Video: youtube.com/watch?v=SNw-zXgv_vA
@@ I Introduction
The last two decades of rapid progress in AI have seen machine learning systems challenge, and even defeat, humans in several tasks (from image recognition to the game of Go, to self-driving). Since the 2019 AlphaPilot AI Drone Innovation Challenge [2, 3], drone racing has emerged as the de facto benchmark for evaluating vision-based aerial autonomy. In 2023, state-of-the-art research has demonstrated that autonomous racing drones can surpass human-level performance in controlled racing arenas [4]. However, the applicability of fast, vision-based autonomous flight to practical commercial operations is still limited, as current systems are typically trained and evaluated in highly controlled (often externally instrumented) environments.
@@
  • Presenting a perception and control stack for autonomous drone racing that does not require fine-tuning with ground truth for residual estimation (and has proven to be resilient to multiple lighting conditions).
  • Releasing pro-level piloting data from the instrumented track using the same format as in [5], adding six new flights by a world champion pilot (Sec. VI-D).
@@
This 2025 Quadcopter Research 25.63m/s
5.04s 21.15m/s
5.60s Human
@@ II Related Work
In car racing, since the 2004/2005 DARPA Grand Challenges, various leagues and research challenges have been established, including the Indy Autonomous Challenge (IAC) and the Abu Dhabi Autonomous Racing League (A2RL)—the latter organizing “Man vs Machine” events for both drone and car racing [6]—aimed at accelerating progress in performance, research, and entertainment.
@@ III Drone Platform
The components listed in this section result in an autonomous drone capable of speeds over 25m/s along an agile indoor trajectory and weighing just 665.5g (without battery).
@@ III-A1 Motors and Propellers
The quadrotor features the T-motor F60 PRO V 2020KV motors, as in [5], paired with HQProp HeadsUp R38 propellers to accommodate the professional pilots’ unanimous preference for a lower blade pitch. This setup results in a Thrust-to-Weight Ratio (TWR) of $\sim$7 at full battery capacity, measured using a thrust bench.
@@ III-A3 Compute
As a companion board, we use the NVIDIA Orin NX module, installed on the A603 carrier board by Seeedstudio. The Orin NX module comes with JetPack 5.1.2 (i.e., Linux Kernel 5.10, Ubuntu 20.04-based root file system, and CUDA 11.4 support). We select the MAXN power mode to maximize core usage and clock frequencies.
@@ III-B1 Autopilot
Our drone requires low-latency, real-time reactive control. This is managed by the FC, running Betaflight (BF) 4.3.2 firmware tuned for the platform [12]. The FC is interfaced using the MultiWii Serial Protocol (MSP), enabling efficient bidirectional data exchange between the FC and the onboard computer. This allows for IMU readings at a rate up to 500Hz over a 1MBaud serial connection (much more than the 10Hz obtainable with the SBUS protocol, as reported in [11]). Additionally, the MSP Override feature provides the onboard computer with complete control over the flight channels. A real-time Linux kernel setup—with priority granted to the MSP process—was crucial to achieve the performance reported in this letter.
@@ III-B2 Vison and Autonomy
Figure 4: Vision stack schematics illustrating the training procedure (top) and the detection pipeline (bottom). The training procedure comprises three steps (designed to minimize human supervision while preserving accuracy): (i) training on the RATM dataset [5]; (ii) auto-labeling, followed by human refinement in CVAT; and (iii) labeling track images with Grounding DINO from human examples. By the final stage, fine-tuning for peak performance in new environments requires as few as 80 human-corrected frames. In the detection pipeline, YOLOv8n detects gate bounding boxes, while MobileNetV3-Small [14] estimates their keypoints. Leveraging GPU acceleration (see Sec. IV-A), the stack achieves per-frame latency under 30 ms on $640\times 640$ images. This yields lower delay and larger input resolution compared to [4], resulting in higher pixel-level precision and potentially improved PnP accuracy.
@@ IV-A Vision Stack
For gate detection, we use You Only Look Once (YOLO), specifically, the YOLOv8n model with $3.2$ million parameters, one class (gates), and input size ($\times640$).
For corner detection, the pixels corresponding to each detected gate region are cropped and forwarded to a custom key-points detection model, MobileNetV3-Small [14] with $1.1$ million parameters. The input size of the model is ($\times256$), and its weights are initially pretrained for a classification task using the ImageNet-1K dataset.
@@ IV-B State Estimation Stack
Our state estimation stack leverages VIO from the Intel T265. Although VIO provides highly accurate short-term estimates, its performance degrades over time, leading to significant drift, especially during fast maneuvers that introduce motion blur. To mitigate this drift, we incorporate detections of the only predefined landmarks in drone racing: the gates. By identifying the corners of the gates in the captured images, the drone’s pose relative to the gate is computed by solving the Perspective-n-Point (PnP) problem, of which only the positional component is used. The global position of the drone is then refined using the gate’s known location in the track layout, whereas the attitude is left uncorrected, as it exhibits a substantially lower drift over time.
Using fixed exposure and carefully tuning the exposure time and gain, we obtain good image brightness with minimal motion blur, ensuring reliable VIO. We enable “mapping”, allowing the device to build and update an internal map for loop closure and small drift correction. In contrast, we disable “relocalization”, which attempts to re-align the estimate after large drifts, and “pose jumping”, which can cause discontinuous estimates. This configuration provided the best balance of VIO accuracy and reliability, while still allowing us to correct larger drifts using gate measurements. The right camera operates at 30Hz, and its frames are synchronized with the corresponding VIO estimates.
@@
The filter state and covariance are initialized to zero, while the process noise is set to $\sigma_{a}^{2}=8$. The filter is then updated whenever a new measurement $\mathbf{z}_{k}$ (a position estimate from a gate detection) is available, using the Kalman filter equations:
@@
in which $K_{k}$ is the Kalman gain, $R$ is the measurement covariance matrix, estimated via Monte-Carlo sampling as described in [4], and $H_{k}$ is the measurement matrix. When several gates are detected in a single camera frame, all relative pose estimates are stacked and processed in the same Kalman filter update step as multiple simultaneous measurements. The resulting drift estimate is constantly used to correct the VIO position estimate. As a final step, the state estimate is refined by fusing it with IMU data from the FC, which operates at 500Hz. This fusion is performed using an Extended Kalman Filter (EKF) with state vector $\mathbf{x}=[\mathbf{q}^{T},\mathbf{p}^{T},\mathbf{v}^{T},\mathbf{b}_{\omega}^{T},\mathbf{b}_{a}^{T}]^{T}\in\mathbb{R}^{16}$ and error state vector $\mathbf{e}=[\mathbf{\psi},\mathbf{\theta},\mathbf{\phi},\mathbf{p}^{T}]^{T}\in\mathbb{R}^{6}$, in which $\mathbf{q}$ represents the drone’s orientation quaternion and $\mathbf{\psi},\mathbf{\theta},\mathbf{\phi}$ its Euler angles representation, $\mathbf{p}$ is its position, $\mathbf{v}$ is its linear velocity, $\mathbf{b}_{\omega}$ is the bias of the gyroscope and $\mathbf{b}_{a}$ is the bias of the accelerometer. The propagation step follows the equations:
@@
Where $\mathbf{P}$ is the state covariance matrix, $\mathbf{F}$ is the state transition matrix, and $\mathbf{W}$ is the Jacobian of the error-state transition model with respect to process noise. The update step is then defined as in 3, 4, 5, but in this case the measurement is $\mathbf{z}_{k}=[\mathbf{\psi},\mathbf{\theta},\mathbf{\phi},\mathbf{p}^{T}]^{T}\in\mathbb{R}^{6}$, representing attitude and gate-corrected position estimates from VIO, and $\mathbf{R}$ is a static diagonal covariance matrix.
@@ IV-C1 Time Optimal Trajectory Generation
Each gate was assigned two waypoints, positioned at the center of the gate along the y- and z-axes. Along the x-axis, the waypoints were placed at -0.4m and +0.4m (-0.4m and +1.25m for the Split-S). This displacement is defined in gate-frame, and it is transformed in world-frame using the known gate yaw angle (in radians) specified in Fig. 5. The world- and gate-reference frame are shown in Fig. 1.
@@ V Racing Tracks
For our experiments, we used two different tracks (Fig. 5). The first track was set up in a flight arena equipped with a motion capture (MoCap) system, providing ground-truth data to analyze speed, gate-to-gate times, traveled distance, state estimation, and trajectory tracking. The second track was prepared in a larger, uninstrumented hangar. In this case, quantitative evaluation was limited to lap and sector times.
@@ V-A2 Motion Capture
The instrumented arena is equipped with a 32-camera Arqus A12 Qualisys MoCap system [^3] https://www.qualisys.com/cameras/arqus, capable of tracking the 6DoF poses of rigid bodies with millimeter-level precision at 275Hz. The drone presented in Sec. III-A was outfitted with five 25mm markers, three on 3D-printed arm extensions, and two near the center of the frame.
@@ V-B1 Technical Specifications
In the larger, uninstrumented hangar, we reproduced, as faithfully as possible, the Track Split-S from [4]. We positioned the gates with the support of a total station, resulting in position errors in the order of a few centimeters. The gates’ orientations had errors of at most ten degrees. We should also note that the bottom gate of the Split-S in our track is 20cm lower compared to the original track due to a different gate interlocking system. The gates’ placement is similar to that of Track RATM. Both tracks feature the Split-S maneuver and similar z-axis variations, although this track has softer curves and lacks the sharp hairpin and spiral segment. The gate poses of the original track are published in the supplementary material of [4].
@@ VI-A Participants
Three professional pilots participated in the experiment, divided into two sessions. The first session included two top-ten-ranked pro-pilots from the Drone Champions League (DCL), namely Thomas Kund (Star23467) and Krutharth M. C. (Ion FPV). The second session featured a world-champion pilot, Minchan Kim (MCK), who had earned multiple titles in recent years—in MultiGP Championship, FAI World Drone Racing Championship (WDRC), and Drone Racing League (DRL)—and led his team to victory in the DCL.
@@ VI-B Runs
The pilots were also allowed to choose how many batteries or hours of practice to use each day, as well as the number of laps to complete with each battery. The autonomous drone always ran one lap per battery instead.
@@ VI-C Performance
The instrumented evaluation allowed the comprehensive analysis of multiple quantitative metrics, including lap times, average speeds, and path lengths. We recorded data for three human pilots and our autonomous system, the latter using both external state estimation (MoCap) as well as fully on-board autonomy (VIO). The results of these flights are presented in Table II and Fig. 6. The autonomous systems (MoCap and VIO) surpassed the humans in average lap time and top speed, although they fell short to one out of the three pilots in terms of best individual results (achieved by MCK). In the instrumented arena, the autonomous system showed a clear reliability advantage, completing all runs without any crash, whereas the human pilots suffered multiple incidents.
@@ VI-D Dataset
Six flights by the fastest pilot, MCK, totaling 240.77s of flying time and 2342.98m of traveled distance, at a top speed of 21.29m/s, are added to our drone-racing-dataset, using the format described in [5].
Figure 7: Comparison of the trajectories followed during the Split-S maneuver by the fastest pilot (MCK) and our autonomous system using only onboard sensing (VIO), sorted by lap time, on the instrumented track.
@@
Figure 9: Lap time distributions and crashes for the fastest pilot and our autonomous system on the uninstrumented track.
@@ VII Uninstrumented Results
The scenario was not entirely unknown, as the basic data required for the vision stack were available. To support the vision module, images of the new location were captured and used to fine-tune the corner detection algorithm. This adjustment allowed for accommodating the significant variations in lighting conditions caused by sunlight streaming through the hangar windows (as tests had to be carried out at all hours of the day). The gate map, essential to support the state estimation module, was created using a total station. Although accurate, this method was considerably less precise than the map generated by the MoCap system.
@@ VII-C Performance
MCK achieved both a lower average and best lap time (see Table II, Fig. 9). The autonomous drone, which was 1.27% faster than MCK on average lap time in the instrumented track, gave up a 3.65% margin to the fastest human. Note that the second and third ranked pro-pilots were $\sim$40-60% slower than MCK in the instrumented track. MCK had 2 crashes, whereas the autonomous system crashed 4 times—3 of which were due to collisions with the human pilot, who was able to recover in two of these cases.